# HOMIE: Humanoid Loco-Manipulation with Isomorphic Exoskeleton Cockpit

### v1

@@ Abstract
Current humanoid teleoperation systems either lack reliable low-level control policies, or struggle to acquire accurate whole-body control commands, making it difficult to teleoperate humanoids for loco-manipulation tasks. To solve these issues, we propose HOMIE, a novel humanoid teleoperation cockpit integrates a humanoid loco-manipulation policy and a low-cost exoskeleton-based hardware system. The policy enables humanoid robots to walk and squat to specific heights while accommodating arbitrary upper-body poses. This is achieved through our novel reinforcement learning-based training framework that incorporates upper-body pose curriculum, height-tracking reward, and symmetry utilization, without relying on any motion priors. Complementing the policy, the hardware system integrates isomorphic exoskeleton arms, a pair of motion-sensing gloves, and a pedal, allowing a single operator to achieve full control of the humanoid robot. Our experiments show our cockpit facilitates more stable, rapid, and precise humanoid loco-manipulation teleoperation, accelerating task completion and eliminating retargeting errors compared to inverse kinematics-based methods. We also validate the effectiveness of the data collected by our cockpit for imitation learning. Our project is fully open-sourced, demos and code can be found in our website.
@@ I Introduction
Humans consistently rely on their physical agility to walk, squat to specific heights, and manipulate objects when performing tasks such as transferring goods in warehouses. While humanoid robots are expected to eventually take over these labor-intensive tasks, achieving full autonomy remains a significant challenge. In the interim, a robust teleoperation system is essential, enabling operators to seamlessly control robots for complex loco-manipulation tasks. These systems not only allow humans to guide robots with precision and fluidity but also serve as a critical tool for collecting high-quality demonstrations to advance autonomous capabilities. Two indispensable components are required to realize this vision: (1) a robust loco-manipulation policy that enables robots to extend operational workspace; (2) a method that allows a single operator to precisely control the robot’s upper-body manipulation and lower-body locomotion simultaneously.
However, most existing efforts to teleoperate humanoid robots focus solely on controlling the upper body [1, 2, 3, 4, 5]. This restriction significantly limits the robot’s operational workspace, as it prevents the robot from moving freely or adjusting its posture (e.g., squatting) to adapt to different task requirements. While some researchers have explored equipping humanoids with loco-manipulation capabilities through reinforcement learning (RL) [6, 7, 8, 9, 10, 11], these approaches often fail to achieve the necessary range of motion, such as squatting to specific heights, which is critical for many real-world tasks. Additionally, many humanoid whole-body teleoperation methods require operators to physically move to control the robot’s motion, making it challenging to maintain precise and stable full-body control over extended distances. Mainstream teleoperation systems typically rely on vision-based methods [1, 3, 4, 5] or heterogeneous exoskeletons [2] to determine end-effector poses, followed by inverse kinematics (IK) to compute joint positions. Nonetheless, both pose estimation and IK solving introduce inaccuracies, which compromises the precision and reliability of teleoperation.
@@
To address these issues, we introduce HOMIE, a novel humanoid teleoperation cockpit composed of a humanoid loco-manipulation policy and an exoskeleton-based hardware system. This cockpit enables a single operator to precisely and efficiently control a humanoid robot’s full-body movements for diverse loco-manipulation tasks. Integrated into simulation environments, our cockpit also enables seamless teleoperation in virtual settings. Specifically, we introduce three core techniques to our RL-based training framework: upper-body pose curriculum, height tracking reward, and symmetry utilization. These components collectively enhance the robot’s physical agility, enabling robust walking, rapid squatting to any required heights, and stable balance maintenance during dynamic upper-body movements, thereby significantly expanding the robot’s operational workspace beyond existing solutions. Unlike previous whole-body control methods that depend on motion priors derived from motion capture (MoCap) data [12], our framework eliminates this dependency, resulting in a more efficient pipeline. Our hardware system features isomorphic exoskeleton arms, a pair of motion-sensing gloves, and a pedal. The pedal design for locomotion command acquisition liberates the operator’s upper body, enabling simultaneous acquisition of upper-body poses. Since the exoskeleton arms are isomorphic to the controlled robot and each glove has 15 degrees of freedom (DoF), which is more than most existing dexterous hands, we can directly set upper-body joint positions from the exoskeleton readings, dispensing with IK and achieving faster and more accurate teleoperation. Moreover, our gloves can be detached from the arms, allowing them to be reused in systems isomorphic to different robots. The total cost of the hardware system is only \$0.5k, which is significantly lower than that of MoCap devices [13].
@@
Through ablation experiments, we validate the effectiveness of each technique in our training framework and demonstrate the robustness of the resulting policies across different robots. Our evaluation shows that the hardware system supports 200% faster and more accurate pose acquisition than previous methods, enabling operators to complete tasks more efficiently than virtual reality (VR)-based approaches. Real-world studies confirm that the trained policies can be deployed directly in the real world, allowing robots to perform diverse loco-manipulation tasks stably in complex environments. We further show that real-world data collected via HOMIE can be effectively used by imitation learning (IL) algorithms, allowing humanoid robots to autonomously execute tasks.
In summary, our core contributions are three-fold:
@@
    We propose HOMIE, the first implementation of a cockpit for humanoid whole-body loco-manipulation teleoperation enables a single operator to fully control the robot and execute diverse tasks seamlessly.
@@
    We first achieve robust humanoid loco-manipulation including squatting under arbitrary continuous changing upper-body poses without using any motion prior.
@@
    Our hardware system supports more precise and rapid whole-body control of humanoids than previous works, reducing task completion times by nearly half.
@@ II Related Works
II-A Dual Arm Robot Teleoperation
@@
Teleoperating dual-arm robots to perform complex manipulation tasks is an efficient way to collect real-world expert demonstration, which can then be used by IL to learn autonomous skills [14, 1, 3, 17, 18]. Some researchers utilize robotic arms identical to the teleoperated ones [19, 14, 15, 16], making joint-matching possible, thus ensuring high accuracy and fast response speed. However, due to the high cost of robotic arms, the establishment of such a system incurs significant expenses. Additionally, teleoperating dexterous hands with these systems is not feasible. An alternative approach is to use VR devices [1, 5] or just a camera [20, 4, 21]. These works use vision-based techniques to capture the operator’s wrist postures and key points of the hands, which are used by IK to calculate the joint positions of the arms and hands. However, due to limitations in the accuracy, inference speed, and difficulty in handling occlusions of pose estimation, such approaches cannot guarantee rapid and accurate pose acquisition. Some researchers try to use MoCap methods [13, 22, 23, 24] to acquire more accurate poses at higher frequencies, but MoCap equipment is very expensive. Moreover, since IK is an iterative method that approximates solutions, even when wrist and hand poses are captured accurately, the limitations of IK may prevent the robot from achieving the desired posture. Another possible solution is an exoskeleton-based teleoperation system, which does not require an additional identical robot, thus the overall cost is relatively low. Some research calculates the end-effector pose of the exoskeleton using Forward Kinematics (FK) and then apply IK to determine the robot’s joint positions, while using computer vision techniques to capture the hand poses [2]. However, these systems are also limited by the inaccuracies of IK and pose estimation. Some studies utilize isomorphic exoskeletons [15, 16], which can also employ joint-matching to teleoperate the robots, ensuring both low cost and high accuracy and control frequency. Neverthless, these systems typically handle robotic arms equipped with grippers, limiting their application to basic manipulation tasks rather than dexterous ones. Since some projects have introduced cheap and reliable motion-sensing gloves [25, 26], redesigning and combining them with an exoskeleton could potentially overcome this limitation, a solution that has not yet been realized in this field. HOMIE is designed to combine all the advantages mentioned above, integrating isomorphic exoskeleton arms with a pair of novel motion-sensing gloves. We will introduce this system in Sec. III-C. A comparison between HOMIE and previous representative teleoperation systems can be found in Tab. I.
II-B Humanoid Whole-body Loco-Manipualtion
@@
To enable robots to perform whole-body loco-manipulation tasks, some researchers focus on model-based optimization algorithms [27, 28, 29, 30, 31], particularly generating locomotion control laws by solving optimal control problems (OCPs). Despite significant efforts to make OCPs computationally tractable, these algorithms still struggle with complex scenarios due to their high computational demands during online processing. Reinforcement Learning (RL)-based algorithms, especially those based on Proximal Policy Optimization (PPO) [32], offer a more powerful alternative. Using these methods, several studies successfully achieve whole-body loco-manipulation in quadruped robots [33, 34, 35, 36], and some teach humanoid robots to traverse various terrains [37, 38, 39, 40, 41, 42, 43, 44] or perform parkour [45]. These achievements motivate researchers to apply the same techniques to humanoid whole-body loco-manipulation [46]. Some studies train whole-body policies for humanoid robots [10], enabling them to act in a manner similar to human operators or even dance with people. Some other research separates the upper and lower body [9, 6, 7, 8, 11], using policies trained by RL to control the lower body while directly setting the joint positions of the upper body, thus helping robots achieve better balance. Despite achieving impressive results, these methods still face several common limitations. First, they often rely on retargeted MoCCap data [12] to get motion prior [47] for training robots. However, obtaining MoCap data is costly, and adapting robots to new poses necessitates additional data collection, which significantly hinders the scalability of these approaches. Second, many of these methods employ vision-based algorithms to estimate the operator’s poses, which lack the precision of exoskeleton-based devices. This limitation reduces the accuracy required for humanoid robots to perform loco-manipulation tasks effectively. Third, these methods generally fail to incorporate the ability to control a robot’s body height. Height control is crucial for handling objects at varying elevations, and its absence severely restricts the robot’s operational workspace. Finally, when issuing locomotion commands, some studies rely on body movement data directly [9, 6, 7, 8], while others use joysticks or pedals [11]. The former approach becomes impractical when operators need to control robots in large-scale environments, whereas the latter offers a more effective solution. However, controlling with joysticks necessitates the use of hands, which may already be occupied by other manual tasks, thereby highlighting the advantage of using pedals for locomotion commands.
@@ III-A System Overview
As shown in Fig. 1, HOMIE consists of a low-level policy $\pi_{loco}$ and an exoskeleton-based hardware system. At any given time $t$, the operator inside the cockpit observes the first point of view (FPV) of the robot through the display. By stepping on the pedal, the operator provides the required locomotion commands $C_{t}=[v_{x,t},\omega_{yaw,t},h_{t}]$ where $v_{x,t}$ is the desired forward or backward speed, $\omega_{yaw,t}$ is the turning speed, and $h_{t}$ is the target height of the robot’s torso. The policy $\pi_{loco}$ controls the robot’s lower-body based on $C_{t}$. Meanwhile, the operator controls the exoskeleton to provide the required joint angles $q_{upper}$ for the robot’s upper-body, which are directly set to the robot. The upper and lower bodies work in coordination, continuously cycling through the process, ultimately enabling teleoperating robots to complete loco-manipulation tasks either in the real world or in simulation. Communications between the cockpit and the robot are achieved via Wi-Fi, allowing operation even when the robot is far from the hardware system. We can collect demonstrations while teleoperating the robot and use them to train an autonomous policy $\pi_{auto}$. Once trained successfully, $\pi_{auto}$ can take over the operator to give $C_{t}$ and $q_{upper}$, thus driving the robot to perform tasks autonomously.
III-B Locomotion Policy
@@ III-B2 Upper-body Pose Curriculum
We use a curriculum learning technique to ensure that $\pi_{loco}$ can still complete locomotion tasks under any continuously varying poses of the robot’s upper-body. We adjust the sampling range of the upper body joint angles using the upper action ratio $r_{a}$. At the start of training, $r_{a}$ is set to 0. Each time the policy drives the robot to track the linear velocity with a reward function that reaches the threshold, $r_{a}$ increases by 0.05, eventually reaching 1. We first sample $r_{a}^{\prime}$ from the probability distribution
$p(x|r_{a})=\frac{20(1-r_{a})\,e^{-20(1-r_{a})x}}{1-e^{-20(1-r_{a})}},\,\,\,r_{a}\in[0,1)$ (2)
and then resample $a_{i}$ by $\mathcal{U}(0,r_{a}^{\prime})$. We actually sample $a_{i}$ by
$a_{i}=\mathcal{U}(0,-\frac{1}{20(1-r_{a})}\ln\left(1-\mathcal{U}(0,1)\left(1-e^{-20(1-r_{a})}\right)\right))$ (3)
As $r_{a}$ increases, the probability distribution gradually transitions from being close to 0 to $\mathcal{U}(0,1)$. This ensures that during the curriculum process, the probability distribution consistently satisfies $p(x|r_{a})>0,\forall x\in[0,1],r_{a}\in[0,1)$. Compared to directly using $\mathcal{U}(0,r_{a})$, this method approaches the final target in a more gradual and smoother manner. For better understanding of Eq. 3, we visualize it in Appendix A. To simulate the continuous changes in upper body movements when controlled by our cockpit, we resample target upper-body poses every 1 second according to the above process. We then use uniform interpolation to ensure that the target movement gradually changes from the current value to the desired value over the 1-second interval. Without this approach, we find that the robot struggles to maintain balance under continuous motions.

### v2

@@ Abstract
Generalizable humanoid loco-manipulation poses significant challenges, requiring coordinated whole-body control and precise, contact-rich object manipulation. To address this, this paper introduces HOMIE, a semi-autonomous teleoperation system that combines a reinforcement learning policy for body control mapped to a pedal, an isomorphic exoskeleton arm for arm control, and motion-sensing gloves for hand control, forming a unified cockpit to freely operate humanoids and establish a data flywheel. The policy incorporates novel designs, including an upper-body pose curriculum, a height-tracking reward, and symmetry utilization. These features enable the system to perform walking and squatting to specific heights while seamlessly adapting to arbitrary upper-body poses. The exoskeleton, by eliminating the reliance on inverse dynamics, delivers faster and more precise arm control. The gloves utilize Hall sensors instead of servos, allowing even compact devices to achieve 15 or more degrees of freedom and freely adapt to any model of dexterous hands. Compared to previous teleoperation systems, HOMIE stands out for its exceptional efficiency, completing tasks in half the time; its expanded working range, allowing users to freely reach high and low areas as well as interact with any objects; and its affordability, with a price of just \$500. The system is fully open-source, demos and code can be found in our website.
@@ I Introduction
Generalizable humanoid loco-manipulation is crucial for integrating humanoid robots into daily life and enabling them to handle labor-intensive tasks. Achieving this requires coordinated whole-body control (WBC) policies that endow robots with both strong athletic capabilities and precise, contact-rich object manipulation skills for interacting with a variety of objects. Teleoperation is a promising technique to realize this vision, leveraging a data-driven approach to create a flywheel effect. However, the field currently faces a significant dichotomy: reinforcement learning (RL)-trained locomotion policies excel at environmental adaptation but lack the interfaces needed for real-time, precise teleoperation [1, 2, 3, 4, 5, 6]. On the other hand, most existing teleoperation systems focus solely on upper-body control without considering the impact of locomotion on the robot’s operational workspace, thereby severely limiting its functionality [7, 8, 9, 10, 11]. This fragmentation creates a lose-lose situation, where robots either sacrifice dexterous manipulation during movement or compromise their workspaces when performing manipulation.
The path forward demands mutual perspective iteration: RL-based training incorporates upper-body teleoperation interfaces without compromising the robot’s athletic ability, and teleoperation system seamlessly integrate locomotion control modules while affording accurate and smooth pose acquisition. In responce, we introduce HOMIE, a semi-autonomous humanoid teleoperation system that integrates a RL policy for body control mapped to a pedal, an isomorphic exoskeleton arm for arm control, and motion-sensing gloves for hand control. This unified cockpit enables a single operator to precisely and efficiently control a humanoid robot’s full-body movements, addressing both humanoid whole-body control and real-time precise teleoperation.
@@
Our RL-based training framework features three core techniques: upper-body pose curriculum for dynamic balance adaptation, height-tracking reward for precise squatting, and symmetry utilization for action regularization and data augmentation. These components collectively enhance the robot’s physical agility, enabling robust walking, rapid squatting to any required heights, and stable balance maintenance during dynamic upper-body movements, thereby significantly expanding the robot’s operational workspace beyond existing solutions and allowing any teleoperation commands to take effect. Unlike previous whole-body control methods that depend on motion priors derived from MoCap data [12], our framework eliminates this dependency, resulting in a more efficient pipeline.
Complementing the training framework, our hardware system features isomorphic exoskeleton arms, a pair of motion-sensing gloves, and a pedal. The pedal design serves as an effective interface for locomotion command acquisition, guiding the robot’s movement while freeing the operator’s upper body. This setup enables simultaneous acquisition of upper-body poses and removes the need for continuous synchronized walking between the operator and the robot. To eliminate the inaccuracies introduced by inverse kinematics (IK) and pose estimation, which are commonly used in mainstream teleoperation systems, we design the exoskeleton arms to be isomorphic to the controlled robot. This allows us to directly set the upper-body joint positions based on the exoskeleton readings, bypassing the need for IK and resulting in faster and more accurate teleoperation. Each of our gloves offers 15 degrees of freedom (DoF), surpassing most existing dexterous hands, enabling them to control a variety of hand types using the same gloves. Additionally, the gloves can be detached from the arms, making them reusable in systems isomorphic to different robots. The total cost of the hardware system is just \$0.5k, significantly more affordable than motion capture (MoCap) devices [13].
@@
Through ablation experiments, we validate the effectiveness of each technique in our training framework and demonstrate the robustness of the resulting policies across different robots. Our evaluation shows that the hardware system supports 200% faster and more accurate pose acquisition than previous methods, enabling operators to complete tasks more efficiently than virtual reality (VR)-based approaches. Real-world studies confirm that the trained policies can be deployed directly in the real world, allowing robots to perform diverse loco-manipulation tasks stably in complex environments. We further show that real-world data collected via HOMIE can be effectively used by imitation learning (IL) algorithms, allowing humanoid robots to autonomously execute tasks. Integrated into simulation environments, our cockpit also enables seamless teleoperation in virtual settings.
In summary, the key contributions of HOMIE are:
@@
    A novel humanoid teleoperation cockpit that combines RL-based loco-manipulation control with an isomorphic exoskeleton and motion-sensing gloves, enabling full-body control by a single operator.
@@
    The first successful implementation of teleoperation-compatible humanoid loco-manipulation, including dynamic squatting, without relying on motion prior data.
@@
    A cost-effective hardware system that supports more precise and faster whole-body control than existing systems, significantly reducing task completion times.
@@ II Related Works
II-A Teleoperation Systems
Teleoperating dual-arm robots to perform complex manipulation tasks is an efficient way to collect real-world expert demonstration, which can then be used by IL to learn autonomous skills [14, 7, 9, 17, 18]. Some researchers utilize robotic arms identical to the teleoperated ones [19, 14, 15, 16], making joint-matching possible, thus ensuring high accuracy and fast response speed. However, due to the high cost of robotic arms, the establishment of such a system incurs significant expenses. Additionally, teleoperating dexterous hands with these systems is not feasible.
@@ II-A Teleoperation Systems
An alternative approach is to use VR devices [7, 11] or just a camera [20, 10, 21]. These works use vision-based techniques to capture the operator’s wrist postures and key points of the hands, which are used by IK to calculate the joint positions of the arms and hands. However, due to limitations in the accuracy, inference speed, and difficulty in handling occlusions of pose estimation, such approaches cannot guarantee rapid and accurate pose acquisition. Some researchers try to use MoCap methods [13, 22, 23, 24] to acquire more accurate poses at higher frequencies, but MoCap equipment is very expensive. Moreover, since IK is an iterative method that approximates solutions, even when wrist and hand poses are captured accurately, the limitations of IK may prevent the robot from achieving the desired posture. Another possible solution is an exoskeleton-based teleoperation system, which does not require an additional identical robot, thus the overall cost is relatively low. Some research calculates the end-effector pose of the exoskeleton using Forward Kinematics (FK) and then apply IK to determine the robot’s joint positions, while using computer vision techniques to capture the hand poses [8]. However, these systems are also limited by the inaccuracies of IK and pose estimation.
Some studies utilize isomorphic exoskeletons [15, 16], which can also employ joint-matching to teleoperate the robots, ensuring both low cost and high accuracy and control frequency. Neverthless, these systems typically handle robotic arms equipped with grippers, limiting their application to basic manipulation tasks rather than dexterous ones. Since some projects have introduced cheap and reliable motion-sensing gloves [25, 26], redesigning and combining them with an exoskeleton could potentially overcome this limitation, a solution that has not yet been realized in this field. HOMIE is designed to combine all the advantages mentioned above, integrating isomorphic exoskeleton arms with a pair of novel motion-sensing gloves. We will introduce this system in Sec. III-C. A comparison between HOMIE and previous representative teleoperation systems can be found in Tab. I.
II-B Whole-body Loco-Manipulation
To enable robots to perform whole-body loco-manipulation tasks, some researchers focus on model-based optimization algorithms [27, 28, 29, 30, 31], particularly generating locomotion control laws by solving optimal control problems (OCPs). Despite significant efforts to make OCPs computationally tractable, these algorithms still struggle with complex scenarios due to their high computational demands during online processing. Reinforcement Learning (RL)-based algorithms, especially those based on Proximal Policy Optimization (PPO) [32], offer a more powerful alternative. Using these methods, several studies successfully achieve whole-body loco-manipulation in quadruped robots [33, 34, 35, 36], and some teach humanoid robots to traverse various terrains [37, 38, 39, 40, 41, 42, 43, 44] or perform parkour [45].
@@ II-B Whole-body Loco-Manipulation
Achievements in quadrupeds motivate researchers to apply the same techniques to humanoid whole-body loco-manipulation [46]. Some studies train whole-body policies for humanoid robots [5], enabling them to act in a manner similar to human operators or even dance with people. Some other research separates the upper and lower body [4, 1, 2, 3, 6], using policies trained by RL to control the lower body while directly setting the joint positions of the upper body, thus helping robots achieve better balance. Despite achieving impressive results, these methods still face several common limitations. First, they often rely on retargeted MoCap data [12] to get motion prior [47] for training robots. However, obtaining MoCap data is costly, and adapting robots to new poses necessitates additional data collection, which significantly hinders the scalability of these approaches. Second, many of these methods employ vision-based algorithms to estimate the operator’s poses, which lack the precision of exoskeleton-based devices. This limitation reduces the accuracy required for humanoid robots to perform loco-manipulation tasks effectively. Third, these methods generally fail to incorporate the ability to control a robot’s body height. Height control is crucial for handling objects at varying elevations, and its absence severely restricts the robot’s operational workspace. Finally, when issuing locomotion commands, some studies rely on body movement data directly [4, 1, 2, 3], while others use joysticks or pedals [6]. The former approach becomes impractical when operators need to control robots in large-scale environments, whereas the latter offers a more effective solution. However, controlling with joysticks necessitates the use of hands, which may already be occupied by other manual tasks, thereby highlighting the advantage of using pedals for locomotion commands.
@@ III-A System Overview
As shown in Fig. 1, HOMIE consists of a low-level policy $\pi_{loco}$ and an exoskeleton-based hardware system. At any given time $t$, the first point of view (FPV) of the robot will be transferred by Wi-Fi to the display inside the cockpit, so the operator can teleoperate the robot with FPV. By stepping on the pedal, the operator provides the required locomotion commands $C_{t}=[v_{x,t},\omega_{yaw,t},h_{t}]$ where $v_{x,t}$ is the desired forward or backward speed, $\omega_{yaw,t}$ is the turning speed, and $h_{t}$ is the target height of the robot’s torso. The policy $\pi_{loco}$ controls the robot’s lower-body based on $C_{t}$. Meanwhile, the operator controls the exoskeleton to provide the required joint angles $q_{upper}$ for the robot’s upper-body, which are directly set to the robot. The upper and lower bodies work in coordination, continuously cycling through the process, ultimately enabling teleoperating robots to complete loco-manipulation tasks either in the real world or in simulation. Communications between the cockpit and the robot are achieved via Wi-Fi, allowing operation even when the robot is far from the hardware system. We can collect demonstrations while teleoperating the robot and use them to train an autonomous policy $\pi_{auto}$. Once trained successfully, $\pi_{auto}$ can take over the operator to give $C_{t}$ and $q_{upper}$, thus driving the robot to perform tasks autonomously.
III-B Humanoid Whole-body Control
@@ III-B2 Upper-body Pose Curriculum
We use a curriculum learning technique to ensure that $\pi_{loco}$ can still complete locomotion tasks under any continuously varying poses of the robot’s upper-body. We adjust the sampling range of the upper body joint angles using the upper action ratio $\rho_{a}$. At the start of training, $r_{a}$ is set to 0. Each time the policy drives the robot to track the linear velocity with a reward function that reaches the threshold, $\rho_{a}$ increases by 0.05, eventually reaching 1. We first sample $\rho_{a}^{\prime}$ from the probability distribution
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p(\rho_{a}^{\prime}|\rho_{a})=\frac{20(1-\rho_{a})\,e^{-20(1-rho_{a})\rho_{a}^{\prime}}}{1-e^{-20(1-\rho_{a})}},\,\,\,\rho_{a}\in[0,1)}$ (2)
and then resample $a_{i}$ by $\mathcal{U}(0,\rho_{a}^{\prime})$. We actually sample $a_{i}$ by
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}a_{i}=\mathcal{U}(0,-\frac{1}{20(1-\rho_{a})}\ln\left(1-\mathcal{U}(0,1)\left(1-e^{-20(1-\rho_{a})}\right)\right))}$ (3)
As $r_{a}$ increases, the probability distribution gradually transitions from being close to 0 to $\mathcal{U}(0,1)$. This ensures that during the curriculum process, the probability distribution consistently satisfies $p(\rho_{a}^{\prime}|\rho_{a})>0,\forall\rho_{a}^{\prime}\in[0,1],\rho_{a}\in[0,1)$. Compared to directly using $\mathcal{U}(0,\rho_{a})$, this method approaches the final target in a more gradual and smoother manner. For better understanding of Eq. 3, we visualize it in Appendix A. To simulate the continuous changes in upper body movements when controlled by our cockpit, we resample target upper-body poses every 1 second according to the above process. We then use uniform interpolation to ensure that the target movement gradually changes from the current value to the desired value over the 1-second interval. Without this approach, we find that the robot struggles to maintain balance under continuous motions.